# Benchmarking digital quantum simulations and optimization above hundreds of qubits using quantum critical dynamics

### v1

@@ Abstract
The real-time simulation of large many-body quantum systems is a formidable task, that may only be achievable with a genuine quantum computational platform. Currently, quantum hardware with a number of qubits sufficient to make classical emulation challenging is available. This condition is necessary for the pursuit of a so-called quantum advantage, but it also makes verifying the results very difficult. In this manuscript, we flip the perspective and utilize known theoretical results about many-body quantum critical dynamics to benchmark quantum hardware and various error mitigation techniques on up to 133 qubits. In particular, we benchmark against known universal scaling laws in the Hamiltonian simulation of a time-dependent transverse field Ising Hamiltonian. Incorporating only basic error mitigation and suppression methods, our study shows coherent control up to a two-qubit gate depth of 28, featuring a maximum of 1396 two-qubit gates, before noise becomes prevalent. These results are transferable to applications such as digitized quantum annealing and match the results of a 133-site optimization, where we identify an optimal working point in terms of both circuit depth and time step.
@@ I Introduction
The quantum Kibble-Zurek mechanism (QKZM) [24, 25, 26] quantifies the relationship between $t_{\mathrm{f}}$ and the number of defects produced during the annealing run. It predicts a universal scaling, a power-law decay, of the density of defects in the final solution as a function of $t_{\mathrm{f}}$ and the system’s critical exponents [27], i.e., a set of numbers characterizing the system’s behavior near its CP. It has been the subject of a range of largely analog experimental studies [28, 29, 30, 31].
We first present an application-oriented benchmarking method that utilizes the predictability and universality of the Kibble-Zurek (KZ) scaling. Those attributes make it ideal for an intuitive and easily scalable metric to assess the quality of large-scale quantum simulations. This is increasingly important as digital quantum computing devices and algorithms have left the infantile stage of a few tens of qubits with error rates prohibiting more than a handful of two-qubit gates [32, 33]. Today’s quantum devices can exceed 100 qubits at error rates that, combined with error mitigation and suppression (EMS) techniques [34], allow a coherent simulation of thousands of two-qubit gates [14, 35].
@@
We go a step further and propose a concrete, application-oriented benchmarking scheme to assess how many two-qubit gate layers of a given application circuit can be coherently simulated given a specific device and, importantly, EMS method. A schematic of the principle is shown in Fig. 1a. The result is not an abstract fidelity but simply the number of circuit layers that can be reliably simulated, which can be directly transferred to other applications.
@@
Next, we apply this benchmark to combinatorial optimization as an application of QA. Although hardware limitations have shifted attention towards the quantum approximate optimization algorithm (QAOA), QA remains an excellent candidate to achieve a computational advantage in quantum optimization [47, 48]. QAOA is an annealing-inspired variational ansatz, with variational layers representing time steps in digitized QA, used in conjunction with a variational quantum eigensolver (VQE) [49]. Recent studies indicate that solving hard optimization problems requires many QAOA layers [50, 51]. At the same time, the optimally converged parameter values of QAOA in the large-layer limit reproduce annealing schedules, meaning QAOA implements digitized QA with variationally found annealing schedules. It is therefore important to better understand digitized QA. An important question is whether the algorithmic error stemming from a finite time step can improve the efficiency, similar to what happens in simulated QA [47]. Here, we identify an optimal working point with respect to the time step and number of circuit layers to minimize the residual energy given the finite hardware resources and how it depends on the system’s minimum energy gap.
@@
In summary, we explore two directions – benchmarking and optimization. Section II provides the necessary technical background to QA, the QKZM, and Trotterized time evolution. In Section III, we introduce and implement our application-oriented benchmark of quantum simulation against universal behavior. Here, we present the main results of this work – a comparison of different levels of EMS on two quantum processors with up to 133 qubits using our QKZM-based benchmarking. Section IV studies digitized QA for solving optimization problems and the effect that the time step has on the quality of the solution. We identify optimal time steps in the presence of different minimum energy gaps on 12 qubits in Section IV.1 and on systems with up to 133 qubits in Section IV.2. We conclude in Section V.
[…]
./figures/abstract_fig_statevector_new.pdf
@@
Figure 1: Application-oriented benchmarking of quantum simulations. a Benchmarking experiments give detailed insight into device characteristics such as error rates and decoherence times. As devices scale, however, it is desirable to have a benchmarking method that resembles the applications that are executed. By simulating critical dynamics and measuring an observable that follows a known universal scaling, we track for how many time steps, which correspond directly to the number of circuit layers, we can reproduce the expected universal scaling. Here, we measure the density of defects produced in quantum annealing, predicted by the Kibble-Zurek scaling. This provides a direct and concrete metric predicting how many layers of two-qubit gates can be coherently simulated given, e.g., a certain device, qubit subset, and error mitigation technique. b For finite system sizes away from the thermodynamic limit, the density of defects exhibits two scaling regimes beyond the Kibble-Zurek scaling that are theoretically known (see main text). These can benchmark deep circuits. c The departure from Kibble-Zurek scaling is determined by the system size and all three scaling regimes can be identified in statevector simulations with $N=10,\ldots,20$ spins (solid lines).
@@ II Digitized quantum annealing and defect production
$H_{\mathrm{M}}=-\sum_{i}\sigma{}^{x}_{i}\ ,\quad H_{\mathrm{P}}=-\sum_{\langle i,j\rangle}J_{ij}\sigma{}^{z}_{i}\sigma{}^{z}_{j}\ ,$ (2)
with Pauli matrices $\sigma{}^{x}_{i}$ and $\sigma{}^{z}_{i}$ acting on site $i$ and couplings $J_{ij}$ between neighboring sites indicated by $\langle i,j\rangle$. The ground state of $H(0)=H_{\mathrm{M}}$ is |$\ket{+}^{N}$. We apply linear schedules $A(s)=(1-s)$ and $B(s)=s$ for simplicity, resulting in
$H(s)=-A(s)\sum_{i}\sigma{}^{x}_{i}-B(s)\sum_{\langle i,j\rangle}J_{ij}\sigma{}^{z}_{i}\sigma{}^{z}_{j}\ .$ (3)
@@
The TFIM is the simplest model exhibiting a QPT [2] in the thermodynamic limit $N\rightarrow\infty$. In the simple case of uniform couplings $J>0$ ($J<0$), the CP at $J_{c}=A/B$ separates the paramagnetic phase ($A\gg JB$) from the doubly-degenerate (anti-)ferromagnetic phase ($A\ll JB$) with all spins (anti-)aligned. In the disordered case, the latter is substituted by a glassy phase, i.e., an energy landscape with many deep local minima [23].
@@ II.1 The Kibble-Zurek mechanism
The QKZM describes the non-equilibrium dynamics of the systems in the realistic setting of crossing this CP with finite annealing time $t_{\mathrm{f}}$, leading to defects in the final solution. Ref. 24 provides an excellent account of the QKZM and the resulting density of defects scaling in a TFIM, both by employing the original (classical) reasoning of the Kibble-Zurek mechanism (KZM) and in terms of a fully quantum description based on Landau-Zener (LZ) theory. Here, a defect refers to the wrong orientation of a spin and the density of defects to the number of defects averaged over all sites. We now summarize the classical description of the KZM following Ref. 24. Classically, the CP is characterized by a diverging correlation length $\xi$ and relaxation time $\tau$. Far away from the CP, $\tau$ is small enough for the system to instantaneously relax to equilibrium. As the system approaches the CP, $\tau$ becomes equal to and, eventually, grows beyond the time scale on which $H$ changes. When this happens, the system can no longer relax to its instantaneous ground state, with reactions to changes slowing down until they halt completely. This is usually referred to as the critical slowdown and subsequent freezing out. Past the CP, as $\tau$ decreases again, the system unfreezes and, crucially, continues evolving from approximately the frozen-out state, resulting in defects in the final state. In summary, the growth of the system’s relaxation time at the CP, which is finite for finite system sizes, determines the necessary rate of change of the system $\mathrm{d}s\,/\mathrm{d}t\,=1/t_{\mathrm{f}}$ required to track the instantaneous ground state. This is the essence of the KZM. It predicts the density of defects as a function of a finite annealing time $t_{\mathrm{f}}$ with a power-law decay [27, 29]
@@
For uniform couplings $J$, defects in the ferromagnetic solution after annealing appear as pairs of spins with opposite sign and can be measured by $\sigma_{i}^{z}\sigma_{j}^{z}$ correlators over all $N_{\mathrm{e}}$ edges of the spin-lattice, i.e., all nearest-neighbor correlators,
@@ II.2 Density of defect-scaling across various regimes
For finite system sizes and slow anneals, i.e., large $t_{\mathrm{f}}$, Ref. 24 describes an exponential drop in defect density following the regime of KZ scaling that is captured by LZ dynamics. In that case, when the anneal is slow enough to never freeze out and to produce less than one defect in the chain on average, the number of defects is proportional to the LZ probability $p_{\mathrm{LZ}}$ of exciting the system,
@@
with $b$ a constant. Following Refs. 52, 53, for adiabatically slow anneals, i.e., even larger $t_{\mathrm{f}}$, the scaling reads
@@
where $a$ is a constant. The expressions of constants $a$ and $b$ can be found in Refs. 24, 53.
@@ III Application-oriented benchmarking through universal scaling
Here, we suggest a direct approach to benchmark close to applications. We simulate the time evolution under a time-dependent Hamiltonian and benchmark the accuracy with which a known universal scaling is replicated. A schematic of our method is shown in Fig. 1a. We propose to measure the density of defects after digitized QA as a function of total annealing time $t_{\mathrm{f}}$. Given a fixed time step, the time $t_{\mathrm{f}}$ directly corresponds to the number of circuit layers in the Trotterized time evolution circuit. The density of defects will decrease according to the predicted scaling up to a certain threshold $t_{\mathrm{f}}$, after which hardware noise dominates. This causes decoherence of the system and an increase of $n_{\mathrm{def}}$, deviating from the predicted KZ scaling. Many other theoretical and experimental studies of the effect of dissipation on the KZ scaling confirm this increase [27, 30, 57]. The number of circuit layers for which the expected scaling of $n_{\mathrm{def}}$ is observed is the number of circuit layers for which coherent simulation was achieved.
@@
This benchmark has several advantages. First and foremost, it yields a concrete, intuitive, and scalable metric; the number of simulable circuit layers as they can be found in countless applications. Even more so since Hamiltonian simulation is a prime application for quantum computers and a building block for many other applications, for example, quantum phase estimation [58] and sampling algorithms [59]. Second, our method can benchmark hardware and EMS algorithms separately or in combination. Third, our method is scalable since no classical verification is required. Fourth, since the KZ scaling is defined in the thermodynamic limit $N\rightarrow\infty$, it is particularly well-suited to benchmark large digital quantum computers without scaling issues. Finally, since finite-size scaling regimes are also well-understood (cf. Section II.2 and Fig. 1b,c), the method is applicable in the setting of scaling to large circuit depths at qubit counts far below the thermodynamic limit.
@@ III.1 Experimental setup
We now discuss benchmarking different levels of EMS on two quantum processors through digitized QA and measuring defect densities. An $N$-spin Ising model with uniform couplings $J=1$ and linear schedules (cf. Section II) is mapped to $N$ qubits. The qubit register is initialized as $\psi(t=0)=\ket{+}^{N}$, the ground state of $H(t=0)=H_{\mathrm{M}}$, and time evolved under the time-dependent Hamiltonian in Eq. 3. The time evolution is implemented using a first-order PF as in Eq. 11 with a time step of $\Delta t$. Throughout this section, we use a time step of $\Delta t=0.5$, which in our experience is sufficiently small to avoid significant algorithmic errors while, at the same time, allowing for larger annealing times. After annealing for a total time $t_{\mathrm{f}}$, we measure the defect density via Eq. 6 to compare to the expected scaling in Eq. 5. Expectation values of observables are estimated from measurements with Qiskit’s Estimator primitive [60].
[…]
[]./figures/kzm_line.pdf
@@
Figure 2: Density of defects scaling in 1D from 100-qubit circuits. Hardware results comparing the density of defects scaling on subsets of 100 qubits on ibm_torino (top) and ibm_sherbrooke (bottom) employing different levels of EMS. Each point corresponds to one additional Trotter layer or time step of $\Delta t=0.5$. Raw simulation results with no EMS (None) are compared to results using combinations of REM, DD error suppression, and pulse-efficient transpilation. Insets show the qubit layout of the respective processor with the chosen qubit subset highlighted. The black dashed line indicates the threshold time $t_{\mathrm{f}}$ until which the simulation stayed true to the expected KZ scaling, shown by the black solid line.
[…]
[]./figures/kzm_heavy_hex.pdf
@@
Figure 3: Density of defects scaling on a heavy-hexagonal lattice from 133- and 127-qubit circuits. a Statevector results comparing the density of defects scaling in a periodic chain, a heavy-hexagonal lattice consisting of two heavy-hex cells ($2\times 1$), and a $7\times 3$ square lattice, all with $N=21$ spins. The heavy-hexagonal lattice exhibits the same $t^{-1/2}$ scaling as predicted for 1D systems, whereas defects in a true 2D lattice show a steeper scaling. b The couplings on the heavy-hexagonal lattice of ibm_torino and ibm_sherbrooke are grouped such that each Trotter layer consists of three layers of $R_{ZZ}$ gates. c Hardware results comparing the density of defects scaling on ibm_torino (left) and ibm_sherbrooke (right) employing different levels of EMS. Each point corresponds to one additional Trotter layer or time step of $\Delta t=0.5$. Raw simulation results with no EMS (None) are compared to results using combinations of REM, DD error suppression, and pulse-efficient transpilation. The black dashed line indicates the time $t_{\mathrm{f}}$ until which the simulation stayed true to the expected KZ scaling, shown by the black solid line.
@@
We employ the IBM Heron processor ibm_torino and the IBM Eagle processor ibm_sherbrooke for our simulations, with 133 and 127 qubits, respectively, and heavy-hexagonal qubit connectivity [61]. Hardware characteristics such as error rates are given in Appendix B. Eagle and Heron processors differ in their native two-qubit gates and how they are physically realized. ibm_sherbrooke’s native two-qubit gate is an echoed cross resonance (ECR) gate implemented via a dispersive coupling mediated by a fixed-frequency resonator [62, 63]. By contrast, ibm_torino’s native two-qubit gate is a CZ gate realized through tunable-frequency couplers. In a tunable coupler architecture, the coupling element is frequency tunable, and driving it can create different two-qubit gates [64, 65]. Most importantly, this results in ibm_torino having a roughly 6×6$\times$6 × shorter two-qubit gate time compared to ibm_sherbrooke and a higher two-qubit gate quality [42].
@@
On both devices, we compare simulations with no EMS to readout error mitigation (REM) [66, 67] as well as to REM combined with dynamical decoupling (DD) [68]. In addition, ibm_sherbrooke allows for pulse-efficient transpilation of our circuits, another method of error suppression [69]. Pulse-efficient transpilation scales hardware native cross-resonance pulses and their echoes (combined, this makes up an ECR gate on ibm_sherbrooke) and is thus not compatible with ibm_torino’s native CZ gates. Appendix A gives a brief technical summary of each method.

### v2

# Benchmarking digital quantum simulations above hundreds of qubits using quantum critical dynamics

@@ Abstract
The real-time simulation of large many-body quantum systems is a formidable task, that may only be achievable with a genuine quantum computational platform. Currently, quantum hardware with a number of qubits sufficient to make classical emulation challenging is available. This condition is necessary for the pursuit of a so-called quantum advantage, but it also makes verifying the results very difficult. In this manuscript, we flip the perspective and utilize known theoretical results about many-body quantum critical dynamics to benchmark quantum hardware and various error mitigation techniques on up to 133 qubits. In particular, we benchmark against known universal scaling laws in the Hamiltonian simulation of a time-dependent transverse field Ising Hamiltonian. Incorporating only basic error mitigation and suppression methods, our study shows reliable control up to a two-qubit gate depth of 28, featuring a maximum of 1396 two-qubit gates, before noise becomes prevalent. These results are transferable to applications such as Hamiltonian simulation, variational algorithms, optimization, or quantum machine learning. We demonstrate this on the example of digitized quantum annealing for optimization and identify an optimal working point in terms of both circuit depth and time step on a 133-site optimization problem.
@@ I Introduction
In a noise-free, closed system, the density of defects is a non-increasing function of $t_{\mathrm{f}}$, and several scaling regimes can be identified as $t_{\mathrm{f}}$ increases [24, 25]. The quantum Kibble-Zurek mechanism (QKZM) [24, 26, 27] quantifies the relationship between $t_{\mathrm{f}}$ and the number of defects produced during the annealing run in a regime of sufficiently slow, yet finite, $t_{\mathrm{f}}$. It predicts a universal scaling, a power-law decay, of the density of defects in the final solution as a function of $t_{\mathrm{f}}$ and the system’s critical exponents [28], i.e., a set of numbers characterizing the system’s behavior near its CP. It has been the subject of a range of largely analog experimental studies [29, 30, 31, 32, 33].
We first present an application-oriented benchmarking method that utilizes the predictability of known universal scaling laws, such as the Kibble-Zurek (KZ) scaling. Those attributes make it ideal for an intuitive and easily scalable metric to assess the quality of large-scale quantum simulations. This is increasingly important as digital quantum computing devices and algorithms have left the infantile stage of a few tens of qubits with error rates prohibiting more than a handful of two-qubit gates [34, 35]. Today’s quantum devices can exceed 100 qubits at error rates that, combined with error mitigation and suppression (EMS) techniques [36], allow a coherent simulation of thousands of two-qubit gates [14, 37].
@@
We go a step further and propose a concrete, application-oriented benchmarking scheme to assess how many two-qubit gate layers of a given application circuit can be reliably simulated given a specific device and, importantly, EMS method. A schematic of the principle is shown in Fig. 1a. The result is not an abstract fidelity but simply the number of circuit layers that can be reliably simulated, which can be directly transferred to other applications. Here, we focus on the KZ scaling, though other universal scalings, e.g. at short times, can also be used [25].
@@
This benchmark is transferrable to applications in time evolution, optimization, quantum machine learning, and variational algorithms. To showcase this, we apply the benchmark to combinatorial optimization specifically as the underlying circuits directly implement digitized QA or the quantum approximate optimization algorithm (QAOA). QAOA is an annealing-inspired variational ansatz, with variational layers representing time steps in digitized QA, used in conjunction with a variational quantum eigensolver (VQE) [49]. Recent studies indicate that solving hard optimization problems requires many QAOA layers [50, 51]. At the same time, the optimally converged parameter values of QAOA in the large-layer limit reproduce annealing schedules, meaning QAOA implements digitized QA with variationally found annealing schedules. It is therefore important to better understand digitized QA. A relevant question is whether the algorithmic error stemming from a finite time step is detrimental or possibly even beneficial, similar to what happens in simulated QA [52]. Here, we identify an optimal working point with respect to the time step and number of circuit layers to minimize the residual energy given the finite hardware resources and how it depends on the system’s minimum energy gap.
@@
In summary, we explore two directions – benchmarking and optimization – and show how the former can be used to guide the design of the latter. Section II provides the necessary technical background to QA, the QKZM, and Trotterized time evolution. In Section III, we introduce and implement our application-oriented benchmark of quantum simulation against universal behavior. Here, we present the main results of this work – a comparison of different levels of EMS on two quantum processors with up to 133 qubits using our QKZM-based benchmarking. Section IV studies digitized QA, using the same underlying circuits as in the benchmarking experiments, for solving optimization problems and showcases how results from Section III are transferable to related applications. We conclude in Section V.
@@
Figure 1: Application-oriented benchmarking of quantum simulations. a Benchmarking experiments give detailed insight into device characteristics such as error rates and decoherence times. As devices scale, however, it is desirable to have a benchmarking method that resembles the applications that are executed. By simulating critical dynamics and measuring an observable that follows a known universal scaling, we track for how many time steps, which correspond directly to the number of circuit layers, we can reproduce the expected universal scaling. Here, we measure the density of defects produced in quantum annealing, predicted by the Kibble-Zurek scaling. This provides a direct and concrete metric predicting how many layers of two-qubit gates can be reliably simulated given, e.g., a certain device, qubit subset, and error mitigation technique. Note that a contains only schematic plots and shows no actual data. b For finite system sizes away from the thermodynamic limit, the density of defects exhibits two scaling regimes beyond the Kibble-Zurek scaling that are theoretically known (see main text). These can benchmark deep circuits. c The departure from Kibble-Zurek scaling is determined by the system size and all three scaling regimes can be identified in statevector simulations with $N=10,\ldots,20$ spins (solid lines). Shown here are results from Trotterized simulations with $\Delta t=0.5$.
@@ II Digitized quantum annealing and defect production
$H_{\mathrm{M}}=-\sum_{i}\sigma{\mathstrut}^{x}_{i}\ ,\quad H_{\mathrm{P}}=-\sum_{\langle i,j\rangle}J_{ij}\sigma{\mathstrut}^{z}_{i}\sigma{\mathstrut}^{z}_{j}\ ,$ (2)
with Pauli matrices $\sigma{\mathstrut}^{x}_{i}$ and $\sigma{\mathstrut}^{z}_{i}$ acting on site $i$ and couplings $J_{ij}$ between neighboring sites indicated by $\langle i,j\rangle$. The ground state of $H(0)=H_{\mathrm{M}}$ is |$\ket{+}^{N}$. We apply linear schedules $A(s)=(1-s)$ and $B(s)=s$ for simplicity, resulting in
$H(s)=-A(s)\sum_{i}\sigma{\mathstrut}^{x}_{i}-B(s)\sum_{\langle i,j\rangle}J_{ij}\sigma{\mathstrut}^{z}_{i}\sigma{\mathstrut}^{z}_{j}\ .$ (3)
@@
The TFIM is the simplest model exhibiting a QPT [2] in the thermodynamic limit $N\rightarrow\infty$. In the simple case of uniform couplings $J>0$ ($J<0$), the CP at $J_{c}=A/B$ separates the paramagnetic phase ($A\gg JB$) from the doubly-degenerate (anti-)ferromagnetic phase ($A\ll JB$) with all spins (anti-)aligned. In the disordered case, the latter is substituted by a glassy phase, i.e., an energy landscape with many deep local minima [53, 54, 23].
@@ II.1 The Kibble-Zurek mechanism
The QKZM describes the non-equilibrium dynamics of the systems in the realistic setting of crossing this CP with finite annealing time $t_{\mathrm{f}}$, leading to defects in the final solution. Ref. 24 provides an excellent account of the QKZM and the resulting density of defects scaling in a TFIM, both by employing the original (classical) reasoning of the Kibble-Zurek mechanism (KZM) and in terms of a fully quantum description based on Landau-Zener (LZ) theory. Here, a defect refers to the wrong orientation of a spin and the density of defects to the number of defects averaged over all sites. In the ferromagnetic phase specifically, defects are domain walls between anti-aligned spins. We now summarize the classical description of the KZM following Ref. 24. Classically, the CP is characterized by a diverging correlation length $\xi$ and relaxation time $\tau$. Far away from the CP, $\tau$ is small enough for the system to instantaneously relax to equilibrium. As the system approaches the CP, $\tau$ becomes equal to and, eventually, grows beyond the time scale on which $H$ changes. When this happens, the system can no longer relax to its instantaneous ground state, with reactions to changes slowing down until they halt completely. This is usually referred to as the critical slowdown and subsequent freezing out. Past the CP, as $\tau$ decreases again, the system unfreezes and, crucially, continues evolving from approximately the frozen-out state, resulting in defects in the final state. In summary, the growth of the system’s relaxation time at the CP, which is finite for finite system sizes, determines the necessary rate of change of the system $\mathrm{d}s\,/\mathrm{d}t\,=1/t_{\mathrm{f}}$ required to track the instantaneous ground state. This is the essence of the KZM. It predicts the density of defects as a function of a finite annealing time $t_{\mathrm{f}}$ with a power-law decay [28, 30]
@@
For uniform couplings $J$, defects in the ferromagnetic solution after annealing appear as domain walls between anti-aligned spins and can be measured by $\sigma_{i}^{z}\sigma_{j}^{z}$ correlators over all $N_{\mathrm{e}}$ edges of the spin-lattice, i.e., all nearest-neighbor correlators. The density of defects is measured as the density of domain walls,
@@ II.2 Density of defect-scaling across various regimes
As mentioned above, the QKZM is not expected to hold for every possible annealing time $t_{\mathrm{f}}$. Deviations from the KZ scaling occur at both very short and very long $t_{\mathrm{f}}$ [25, 55, 56]. For finite system sizes and very slow anneals, i.e., large $t_{\mathrm{f}}$, Ref. 24 describes an exponential drop in defect density following the regime of KZ scaling that is captured by LZ dynamics. In that case, when the anneal is slow enough to never freeze out and to produce less than one defect in the chain on average, the number of defects is proportional to the LZ probability $p_{\mathrm{LZ}}$ of exciting the system,
@@
with $b$ a constant. Following Refs. 57, 58, for adiabatically slow anneals, i.e., even larger $t_{\mathrm{f}}$, the scaling reads
@@
where $a$ is a constant that depends on the derivatives of the annealing schedules [57]. The expressions of constants $a$ and $b$ can be found in Refs. 24, 58.
@@ III Application-oriented benchmarking through universal scaling
Here, we suggest a direct approach to benchmark close to applications. We simulate the time evolution under a time-dependent Hamiltonian and benchmark the accuracy with which a known universal scaling is replicated. A schematic of our method is shown in Fig. 1a. We propose to measure, for example, the density of defects after digitized QA as a function of total annealing time $t_{\mathrm{f}}$. Given a fixed time step, the time $t_{\mathrm{f}}$ directly corresponds to the number of circuit layers in the Trotterized time evolution circuit. In the KZ regime, the density of defects will decrease according to the predicted scaling up to a certain threshold $t_{\mathrm{f}}$, after which hardware noise becomes dominant and leads to a deviation from the KZ scaling. The number of circuit layers for which the expected scaling of $n_{\mathrm{def}}$ is observed is the number of circuit layers for which reliable simulation was achieved.
In particular, decoherence of the system will lead to a deviation of the defect density from the predicted KZ scaling. This has been previously confirmed in several numerical and analog experiments that studied the effect of dissipation on the KZ scaling [62, 63, 28, 31]. These studies show that different kinds of noise can cause decoherence of the system. However, the behavior of the density of defects as noise accumulates, and whether it increases or decreases, depends on the noise.
In this regard, it is important to note that observing a $t_{\mathrm{f}}^{-1/2}$ scaling may not be a sufficient condition to claim coherent quantum dynamics. For instance, a purely classical diffusion model may reproduce the same density of defects scaling [64, 31]. This is particularly important for analog quantum annealers. In the presence of sufficiently strong thermal noise that destroys coherence, the density of defects could still decrease with $t_{\mathrm{f}}$, since the machine may behave like a classical thermal annealer (at low temperature) or as a machine featuring incoherent quantum tunneling events [65, 66, 67]. These processes could still lower the density of defects. Indeed, this effect has been observed in Ref. 31, necessitating additional analysis to reasonably prove the existence of coherent annealing [68].
In our digital quantum setting, the situation is different: the presence of hardware noise never yields a behavior that resembles a classical thermal limit. To prove this, in Appendix A, we study the impact of simulated hardware noise with varying strength. We observe that introducing hardware noise qualitatively changes $n_{\mathrm{def}}$ as a function of $t_{\mathrm{f}}$. Even for large $t_{\mathrm{f}}$ and accumulating noise, we do not recover a decreasing $n_{\mathrm{def}}(t_{\mathrm{f}})$, contrary to the noisy analog quantum annealing case [31]. Therefore, the density of defects following the expected scaling constitutes evidence of a noise-free digital simulation up to a threshold $t_{\mathrm{f}}$. This argument is further corroborated by the fact that the noise model is in very good agreement with hardware results (Appendix A).
@@
The proposed benchmark has several advantages. First and foremost, it yields a concrete, intuitive, and scalable metric; the number of simulable circuit layers as they can be found in countless applications. Even more so since Hamiltonian simulation is a prime application for quantum computers and a building block for many other applications, for example, quantum phase estimation [69] and sampling algorithms [70]. Second, our method can benchmark hardware and EMS algorithms separately or in combination. Third, our method is scalable since no classical verification is required. Fourth, since the KZ scaling is defined in the thermodynamic limit $N\rightarrow\infty$, it is particularly well-suited to benchmark large digital quantum computers without scaling issues. Finally, since finite-size scaling regimes are also well-understood (cf. Section II.2 and Fig. 1b,c), the method is applicable in the setting of scaling to large circuit depths at qubit counts far below the thermodynamic limit.
@@ III.1 Experimental setup
We now discuss benchmarking different levels of EMS on two quantum processors through digitized QA and measuring defect densities. An $N$-spin Ising model with uniform couplings $J=1$ and linear schedules (cf. Section II) is mapped to $N$ qubits. The qubit register is initialized as $\psi(t=0)=\ket{+}^{N}$, the ground state of $H(t=0)=H_{\mathrm{M}}$, and time evolved under the time-dependent Hamiltonian in Eq. 3. The time evolution is implemented using a first-order PF as in Eq. 11 with a time step of $\Delta t$. Throughout this section, we use a time step of $\Delta t=0.5$. We justify the choice of this time step in Appendix B, showing results of noiseless statevector simulations with different time steps and comparing them to continuum results. Although smaller time steps would yield higher accuracy, more Trotter steps, i.e., deeper circuits, would be required to explore the relevant KZ regime. Our results show that $\Delta t=0.5$ is a reasonable compromise to avoid significant algorithmic errors and allow for sufficiently large annealing times. Hardware improvements will eventually allow for smaller time steps and thereby increase the overall accuracy of the results, particularly at short time scales. After annealing for a total time $t_{\mathrm{f}}$, we measure the defect density via Eq. 6 to compare to the expected scaling in Eq. 5. Expectation values of observables are estimated from measurements with Qiskit’s Estimator primitive [71].
@@
Figure 2: Density of defects scaling in 1D from 100-qubit circuits. Hardware results comparing the density of defects scaling on subsets of 100 qubits on ibm_torino (top) and ibm_sherbrooke (bottom) employing different levels of EMS. Each point corresponds to one additional Trotter layer or time step of $\Delta t=0.5$. Raw simulation results with no EMS (None) are compared to results using combinations of REM, DD error suppression, and pulse-efficient transpilation. Insets show the qubit layout of the respective processor with the chosen qubit subset highlighted. The grey shading indicates decoherence and deviation from the expected KZ scaling shown by the black solid line.
@@
Figure 3: Density of defects scaling on a heavy-hexagonal lattice from 133- and 127-qubit circuits. a Statevector results comparing the density of defects scaling in a periodic chain, a heavy-hexagonal lattice consisting of two heavy-hex cells ($2\times 1$), and a $7\times 3$ square lattice, all with $N=21$ spins. The heavy-hexagonal lattice exhibits the same $t^{-1/2}$ scaling as predicted for 1D systems, whereas defects in a true 2D lattice show a steeper scaling. b The couplings on the heavy-hexagonal lattice of ibm_torino and ibm_sherbrooke are grouped such that each Trotter layer consists of three layers of $R_{ZZ}$ gates. c Hardware results comparing the density of defects scaling on ibm_torino (left) and ibm_sherbrooke (right) employing different levels of EMS. Each point corresponds to one additional Trotter layer or time step of $\Delta t=0.5$. Raw simulation results with no EMS (None) are compared to results using combinations of REM, DD error suppression, and pulse-efficient transpilation. The grey shading indicates decoherence and deviation from the expected KZ scaling shown by the black solid line.
@@
We employ the IBM Heron processor ibm_torino and the IBM Eagle processor ibm_sherbrooke for our simulations, with 133 and 127 qubits, respectively, and heavy-hexagonal qubit connectivity [72]. Hardware characteristics such as error rates are given in Appendix D. Eagle and Heron processors differ in their native two-qubit gates and how they are physically realized. ibm_sherbrooke’s native two-qubit gate is an echoed cross resonance (ECR) gate implemented via a dispersive coupling mediated by a fixed-frequency resonator [73, 74]. By contrast, ibm_torino’s native two-qubit gate is a CZ gate realized through tunable-frequency couplers. In a tunable coupler architecture, the coupling element is frequency tunable, and driving it can create different two-qubit gates [75, 76]. Most importantly, this results in ibm_torino having a roughly 6×6$\times$6 × shorter two-qubit gate time compared to ibm_sherbrooke and a higher two-qubit gate quality [44].
@@
On both devices, we compare simulations with no EMS to readout error mitigation (REM) [77, 78] as well as to REM combined with dynamical decoupling (DD) [79]. In addition, ibm_sherbrooke allows for pulse-efficient transpilation of our circuits, another method of error suppression [80]. Pulse-efficient transpilation scales hardware native cross-resonance pulses and their echoes (combined, this makes up an ECR gate on ibm_sherbrooke) and is thus not compatible with ibm_torino’s native CZ gates. Appendix C gives a brief technical summary of each method.